How long will it take a given sum of money to increase 4 times it's present value when compounded half yearly at 7% rate of interest?


1
Expert's answer
2020-10-02T06:37:20-0400

20.149 years

calculation:

"A=P(1+\cfrac{r}{n})^{nt}"


A=4P

 r=0.07

 n=2

t is time taken

"4P=P(1+\cfrac{0.07}{2})^{2t}"


"\cfrac{4P}{P}=\cfrac{P}{P}(1+0.035)^{2t}"


"4=(1.035)^{2t}"


"4=(1.035^2)^t"


"t=\cfrac{log 4}{log 1.035^2}"


"t=20.14879168"